Q1: Find Vc(t) for 20 using general solution + 10V 2011 0.15F 2

Help me know please , I have an exam its time 20 minutes
Answer #1

Given Commected to Jov source at tuo Switch is and Capacitor ºs opeos, voltage asos Capacitor is Kro) 22 10V. 7 5602 VCO) 20v

T = Req c. Time Constant 232 Req 52 Req Req= 542 5+2 olto 2 To P= +0.15 1-5 Sec kle that Vect )= Voca) + ( vecot) - veca) Jet